# Conventional and practical metallic superconductivity arising from repulsive Coulomb coupling

### v1

@@ IV.3 Application to the three dimensional electron gas
that allows the existence of superconductivity. This is analogous to the fact that the Eliashberg coupling $\lambda$ [40] must exceed the Coulomb pseudopotential $\mu^{*}$ to generate superconductivity. It should be noted that, the result of absence of superconductivity is technically based on a lower-bound on $T_{c}$, which leaves open the possibility of weak superconductivity. The parameters $U_{0,1}$ can be calculated for the electron gas using Eq. 35. As shown in Fig. 6, the electron gas can satisfy the above superconductivity condition for only one of the values of $r_{s}=2.8$ considered for $\omega_{c}=E_{F}$. Note that this SC for $r_{s}=2.8$ disappears for $\omega_{c}<0.9E_{F}$, since reducing $\omega_{c}$ moves each point to lower $U_{1}\omega_{c}$. One caveat for these results is that unlike the case of SC induced by electron-phonon coupling [39], the electron gas points depend on the width $k_{c}=10^{-3}k_{F}$ of the momentum-space shell over which the order parameter $\phi_{m}(k)$ is uniform. In fact, smaller $k_{c}$ turns out to be more favorable to superconductivity so that for $k_{c}=10^{-2}k_{F}$ would render the range of $r_{s}$ non-superconducting while $k_{c}=10^{-6}k_{F}$ would lead to most of the range in Fig. 6 superconducting. However, the value of $k_{c}$ is likely to limit the gap $\Delta$ once one goes beyond the linearized gap equation through the relation $k_{c}>\Delta/v_{F}$. Therefore, for practical SC, one should limit to $k_{c}>10^{-4}k_{F}$. Even for parameters with superconductivity, the $T_{c}$ is expected to drop for small $\omega_{c}$, which can be checked by explicit numerical solution of the gap equation Eq. 26. It is however clear from the results in Fig. 6 that $T_{c}$ is likely to be small or absent in much of the parameter regime of the 3D electron gas.
@@ V Summary and Conclusion
where $\omega_{p}$ is the plasma frequency. We note that Eqs. (41) and (42) here, while being the leading order results in $q$ and $\omega$ respectively, refer to the limits $qv_{F}\gg\omega$ and $qv_{F}\ll\omega$, respectively. We note that Eqs. (41) and (42) respectively indicate the low frequency screening and the high frequency anti-screening behaviors, with the screened interaction $u(q,\omega)=v_{c}(q)/\epsilon(q,\omega)$ behaving as $u<v_{c}$ and $u>v$ for $\omega\ll qv_{F}$ and $\omega\gg qv_{F}$. Note also that for $\omega\gg\omega_{p}$, the Coulomb interaction is unscreened. All of these results are exact, and do not lead to any change in the sign of the effective or screened interaction, and this cannot lead to an effective attraction arising from electron-electron interaction. We note the crucial feature that, although both the high- (Eq. (42)) and low- (Eq. (41)) effective interaction remains repulsive, there is an intermediate frequency regime defined by $qv_{F}<\omega<\omega_{p}$, where the effective dynamically screened interaction becomes attractive since in this regime of $\omega<\omega_{p}$, the dielectric screening in Eq. (42) is negative. This emergent attraction between electrons is entirely a many-body retardation effect, which could, in principle, lead to pairing. The actual attractive region, as described in Sec. II, is larger than the asymptotic considerations of Eqs. (41) and (42) indicate.
@@
In Secs. III and IV, we provide the results for $T_{c}$ based on the minimal BCS theory and the more complete Migdal-Eliashberg theory, respectively. The naive BCS theory predicts absurdly high $T_{c}$ ($\sim 100$$1000$ K) for normal metals arising from plasmon-induced pairing, with the fundamental reason behind this absurdity being the very high electronic energy scales of metals ($\sim 10^{4}$ K) and the basic electron-plasmon coupling strength, $r_{s}$, also being large, $r_{s}\sim 5$ for metals. There are two related reasons for the manifest failure of the BCS theory here: the presence of the underlying omnipresent Coulomb repulsion (i.e. the so-called $\mu^{*}$ effect) suppressing $T_{c}$ and the vertex corrections. Neither of these effects is included in the BCS theory, making its predictions unreliable. An ad hoc inclusion of a $\mu^{*}$ could certainly push $T_{c}$ down, and in fact it is easy to suppress $T_{c}$ to zero simply by positing that $\mu^{*}$ is very large in this problem—e.g. a $\mu^{*}>1$ will endure that normal metals do not undergo plasmon induced superconductivity, but a hypothetical system with very low density (and very large $r_{s}$) still might since such a large $r_{s}$ enables overcoming the repulsive $\mu^{*}$ effect. This is, however, not doing theoretical physics, it is just data fitting, since there is no reliable method for calculating $\mu^{*}$ quantitatively. The problem of vertex corrections is conceptually even greater. The quantitative success of the theory of superconductivity for electron-phonon interaction is based entirely on the existence of a small parameter (even when the basic coupling is strong, as, e.g., for Pb where the dimensionless electron-phonon coupling $\sim 1.7$) which is essentially the ratio of the bosonic energy scale to the fermionic energy scale (or their group velocity ratio). For phonons, this ratio is very small since the phonon Debye energy $\sim 10^{2}$ K and the electron Fermi energy $\sim 10^{4}$ K. This is the celebrated Migdal theorem which ensures that vertex corrections are negligible, leading to the Migdal-Eliashberg theory. There is no Migdal theorem whatsoever for the electron-electron interaction induced superconductivity by virtue of the fact that both the bosonic glue (e.g. plasmons) and the electrons, by definition, have the same energy scales, and hence the energy (or velocity) scale difference, which is the key to the Migdal theorem, does not exist for this problem by definition. The nonexistence of the Migdal theorem makes any calculation of $T_{c}$ inherently unreliable for the current problem. In Sec. IV, we elaborate on this matter by providing a detailed analysis of the Migdal-Eliashberg theory as applied to 3D normal metals in the context of the electron interaction induced superconductivity. The application of Eliashberg theory, because of Migdall’s theorem, requires us to identify a frequency cut-off $\omega_{c}$, where the Eliashberg gap equation is solved. Additionally, the long-ranged nature of the Coulomb interaction requires us to consider a finite momentum space cutoff as well. Choosing reasonable values of the cut-off, we find that the electron-gas to be non-superconducting in much of the range of $r_{s}$. We therefore expect $T_{c}$ to be small even for the values of $r_{s}\sim 3$ where superconductivity is found for $\omega_{c}=E_{F}$. Corrections to RPA, which is the basis of our interaction model, are likely to become more important at higher $r_{s}$. The range of $r_{s}$ over which the gas is superconducting increases with lowering $k_{c}$. However, a $k_{c}$ much smaller than $10^{-3}k_{F}$ likely corresponds to a sub-kelvin gap superconductor. Furthermore, the superconducting gap or even the presence of superconductivity is found to be drastically reduced by lowering $\omega_{c}$, which is likely required by Migdall’s theorem to avoid significant vertex correction effects. Ultimately, $\omega_{c}$ needs to be chosen to be small to avoid vertex corrections but too low a value of $\omega_{c}$ would under-estimate $T_{c}$. Our conclusion is that in all likelihood any plasmon induced $s$-wave metallic superconductivity has vanishing $T_{c}$. In fact, we find that even the $T=0$ system is not ordered in many cases, i.e., electron-electron interaction simply does not lead to $s$-wave superconductivity in 3D normal metals. It is still however possible that the approximations in our analysis of the Eliashberg equation miss a SC state with a very low $T_{c}$. We mention that our work does not rule out an exponentially low $T_{c}$ for normal metals in higher angular momentum channels arising from the Kohn-Luttinger mechanism, which we do not consider in the current work.
@@
One could legitimately ask whether our conclusion of there being no plasmon-induced superconductivity applies to systems other than 3D jellium electron liquids, which is the explicit case we consider. We believe that our work should apply to all situations where a jellium electron gas model applies since there is nothing specific to a 3D model of metals utilized in our work. The fact that energy scales cannot be separated, and hence vertex corrections and Coulomb repulsion effects are important, but cannot be included in the BCS-Migdal-Eliashberg theories, applies to all situations, including much-studied 2D electron liquids, where the effective $r_{s}$ is often larger (and the effective $\omega_{p}$ smaller) because of 2D electron liquids, mostly being doped semiconductors, are inherently low-density metals. The fact that 2D plasmon dispersion goes as $q^{1/2}$ and thus vanishes at long wavelength is a matter of profound inconsequence in our theoretical considerations since all the results involve some type of frequency-momentum integrations, and what matters is the plasmon energy at $k_{F}$ (or the plasmon energy at the momentum where it enters the electron-hole continuum). In fact, the 2D theory for plasmon-induced metallic superconductivity is identical to our 3D theory, with exactly the same conclusions. There have been claims [42] that having a bilayer 2D systems with acoustic plasmons [43] may facilitate plasmon-induced superconductivity, perhaps motivated by the fact that acoustic plasmons have the same linear-in-momentum energy dispersion as acoustic phonons. But this claim is misguided and misleading because the Migdal theorem still does not apply, and the standard SC theories (i.e. BCS-Migdal-Eliashberg) would produce results very similar to what we get in our work, again with unreliable and unreasonably high $T_{c}$. The problem of the inapplicability of the Migdal theorem does not become any less stringent just by virtue of having acoustic plasmons in the system since the bosons and the fermions are all arising from exactly the same Hamiltonian.
@@
We mention that the manifest failure of the Migdal theorem for a pairing mechanism does not immediately lead to the conclusion that superconductivity may not occur due to that mechanism. All it implies is the failure of our existing theories, the BCS-Migdal-Eliashberg theory, to describe any possible superconductivity arising from that mechanism. Thus, the absurdity of blindly applying the BCS theory for plasmon induced pairing in 3D metals leading to a predicted $T_{c}\sim 10^{2}$$10^{3}$ K only implies, by itself, that such a theory is not useful in discussing superconductivity in this context. For example, if the problem could be solved exactly in some magical manner, where the issue of the inapplicability of the Migdal theorem is no longer relevant, we cannot rule out the existence of a plasmon-induced $T_{c}$. The work in our Sec. IV indicates such a scenario to be extremely unlikely, but we have not solved the problem exactly. We do believe that if the plasmon mechanism, in an unlikely (but not impossible) scenario, leads to superconductivity in normal metals, the associated $T_{c}$ is likely to be impractically small by virtue of the fact that the direct effect of Coulomb repulsion will always be a huge challenge, i.e., the effective mu* is likely to be large negating any attractive pairing trend induced by plasmons. The fact that the bare Coulomb interaction between electrons is repulsive added to the fact that plasmons are intrinsic electronic phenomena make it essentially impossible for the plasmon mechanism to lead to any ‘practical’ cinventional superconductivity in our opinion.
@@
We now comment on the possibility of electron interaction induced superconductivity in strictly lattice strongly correlated Hamiltonians described by a tight binding Hamiltonian with mostly local interactions. The paradigmatic model is the Hubbard model with onsite interaction and nearest neighbor hopping. This model, which is complementary to our jellium based free electron Hamiltonian with long range Coulomb interaction, has been much studied in the literature motivated by high-$T_{c}$ cuprate superconductors. Our work does not directly apply to this model although the inapplicability of the Migdal theorem as well as the $\mu^{*}$ problem exist here too. The question of superconductivity in the repulsive Hubbard model is an extensively studied problem, which is well beyond the scope of the current work. We mention two relevant works. In Ref. [4], the authors conclude that for asymptotically weak interaction, where the weak coupling perturbative renormalization group should apply, there is a $T_{c}$ which is exponentially small. This is akin to the celebrated Kohn-Luttinger superconductivity in electron liquids. The authors speculate that if their perturbative results can be uncritically extended to larger interaction strength, then a $T_{c}$ describing ‘regular’ superconductivity (i.e. not exponentially small) may emerge. But, this conclusion has been directly refuted in a later quantum Monte Carlo and DMRG calculations which assertively claims the absence of superconductivity in the repulsive Hubbard model for any interaction strength [44]. But a later numerical work found that a slight modification of the tight binding kinetic energy term to include the next-nearest-neighbor hopping in a generalized Hubbard model does manifest superconductivity, but most likely with rather low $T_{c}$, and in higher angular momentum channel (e.g. $d$-wave) [45]. We believe that the superconductivity discussed in the context of the Hubbard model in these (and possibly other) publications are essentially low-temperature superconductivity in higher orbital angular momentum, whose physics is connected to Kohn-Luttinger superconductivity in the electron liquid context, and not the ‘conventional’ superconductivity with ’practical’ transition temperature of interest in the current work. Certainly, the superconductivity discussed in the strongly correlated lattice models of electron interactions is not plasmon-induced in the sense we describe in the current work, and is therefore outside the scope of the current work. In this context, there are recent theories of superconductivity in various graphene and TMD based 2D layers with rather flatbands (which are very far from our electron liquid metals in the current work), where the superconductivity (with low $T_{c}$) may very well be arising from electron-electron interactions, but again such superconductivity is most likely a variation on the Kohn-Luttinger mechanism and is not generated by the attractive pairing glue of the virtual plasmon exchange between the electrons [46, 5, 47].

### v2

@@ IV.3 Application to the three dimensional electron gas
that allows the existence of superconductivity. This is analogous to the fact that the Eliashberg coupling $\lambda$ [mcmillan1968transition] must exceed the Coulomb pseudopotential $\mu^{*}$ to generate superconductivity. It should be noted that, the result of absence of superconductivity is technically based on a lower-bound on $T_{c}$, which leaves open the possibility of weak superconductivity with very low $T_{c}$. The parameters $U_{0,1}$ can be calculated for the electron gas using Eq. 37. As shown in Fig. 9, the electron gas can satisfy the above superconductivity condition for only one of the values of $r_{s}=2.8$ considered for $\omega_{c}=E_{F}$. Note that this SC for $r_{s}=2.8$ disappears for $\omega_{c}<0.9E_{F}$, since reducing $\omega_{c}$ moves each point to lower $U_{1}\omega_{c}$. One caveat for these results is that unlike the case of SC induced by electron-phonon coupling [marsiglio2008electron], the electron gas points depend on the width $k_{c}=10^{-3}k_{F}$ of the momentum-space shell over which the order parameter $\phi_{m}(k)$ is uniform. In fact, smaller $k_{c}$ turns out to be more favorable to superconductivity so that for $k_{c}=10^{-2}k_{F}$ would render the range of $r_{s}$ non-superconducting while $k_{c}=10^{-6}k_{F}$ would lead to most of the range in Fig. 9 superconducting. However, this is an artifact of the momentum space singularities in the $T=0$ polarizability function $\varpi$ that would be rounded at finite temperature $T$. This limits the validity of our results to the regime $k_{c}>T_{c}/v_{F}$. Therefore, for practical SC, one should limit to $k_{c}>10^{-4}k_{F}$. Even for parameters with superconductivity, the $T_{c}$ is expected to drop for small $\omega_{c}$, which can be checked by explicit numerical solution of the gap equation Eq. 28. However, it is clear from the results in Fig. 9 that $T_{c}$ is likely to be small or absent in much of the parameter regime of the 3D electron gas.
IV.4 Application to the two dimensional electron gas
The analysis for the superconductivity in the two dimensional electron gas parallels the analysis in the previous subsection except that the dimensionless Lindhard polarizability is replaced by the corresponding function for the two dimensional electron gas [Stern1967]:
$\varpi(q,\omega)=-(2z)^{-1}[2z-f(z-iu)-f(z+iu)],$ (43)
where $f(w)=w\sqrt{1-w^{-2}}$ and the square root is defined to choose the branch with the positive real part. The dimensionless Coulomb interaction in the two dimensional case is written as $U(q,\infty)=N(0)V_{c}(q)=r_{s}/z\sqrt{2}$, $r_{s}$ is the dimensionless 2D electron gas parameter defined in the last section.
By repeating the calculations in the last sub-section with the above interaction and polarizability, we can calculate the critical value of $U_{0}$ as a function of $U_{1}\omega_{c}$ for the 2D electron gas at various $r_{s}$. The results of this calculation are demarcated by the square symbols in Fig. 9. The 2D electron gas seems to be superconducting at a lower value of $r_{s}$ relative to the electron gas because of which we have plotted the results for $\omega_{c}=0.5E_{F}$ instead of $\omega_{c}=E_{F}$ used for the 3D electron gas. One expects the relevant value of $\omega_{c}\ll E_{F}$ to be much smaller than $E_{F}$, so we do not expect the electron gas to be superconducting in this range of $r_{s}$ for either two dimensional or three dimensional electron gas. Also note that the cutoff factor $k_{c}$ introduced to account for the finite temperature effect on the polarization function $\varpi$ has a stronger $k_{c}/k_{F}$ correction in $2D$ as opposed to the $log(k_{F}/k_{c})$ correction in 3D. This should be interpreted as the thermal corrections to RPA that are ignored in this work being a larger correction in 2D.
Figure 10: Critical temperature $T_{c}$ as a function of the cutoff scales $k_{c}$ and $\omega_{c}$ for the electron gas with different parameters. The panels (a)-(d) respectively contain results for (a) 3D electron gas at $r_{s}=1$; (b) 3D electron gas at $r_{s}=5$; (c) 2D electron gas at $r_{s}=1$ and (d) 2D electron gas at $r_{s}=5$. The color scale represents $\log_{10}(T_{c}/E_{F})$, and the black line indicates the boundary $k_{c}=T_{c}/v_{F}$, above which the results become unphysical. These results show the electron gas to be superconducting within RPA at a $T_{c}$ depending on $r_{s}$ and dimensions. However, the $T_{c}$ becomes negligible for $\omega_{c}\ll E_{F}$.
IV.5 Results for $T_{c}$ in various regimes
The previous subsections discussed conditions for the 2D and 3D electron gas to support a finite $T_{c}$ when the frequency dependence of the Fermi surface averaged screened Coulomb interaction $U(\omega)$ could be approximated to be linear (Eq. 40). In addition, the pairing potential $\phi_{m}$ was assumed to have a BCS-like ansatz. Although these approximations are reliable in regimes with low $T_{c}$, they are still uncontrolled approximations. The main motivation for these approximations is that the Eliashberg equations themselves, i.e. Eq. 28, are notably difficult to solve when the transition temperature is small. Furthermore, these approximations make connections to the standard Allen-Dynes [Allen1975] and Coulomb pseudopotntial [morel1962calculation] formalism.
In this sub-section, we present direct solutions to the Eliashberg equation Eq. 28, where $U(\omega)$ is the Fermi-surface averaged screened Coulomb interaction for the 3D and 2D electron gases discussed in the previous two sub-sections. The transition temperature $T_{c}$ is determined by finding when a variational estimate of the lowest eigenvalue of Eq. 29 drops below 1-1. Two complementary numerical approaches are employed depending on the relative scale of $T_{c}$ and the cutoff $\omega_{c}$. In the regime $T_{c}\sim\omega_{c}$, the dimension of the matrix $U(\omega_{m}-\omega_{m^{\prime}})$ is small, allowing the Eliashberg equation to be solved directly by computing its lowest eigenvalue. However, this regime typically fails to satisfy the physically relevant cutoff or limitation conditions $k_{c}\ll T_{c}/v_{F}$ and $\omega_{c}\ll E_{F}$, which will be discussed later, and therefore is not generally applicable.
In the regime $T_{c}\ll\omega_{c}$, the Eliashberg equation Eq. 29 can be approximated by the integral eigenvalue problem
$\psi(\omega)=-\frac{1}{2}\int_{\pi T}^{\omega_{c}}d\omega^{\prime}\,\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{\sqrt{\omega\omega^{\prime}}}\psi(\omega^{\prime}).$ (44)
Direct evaluation of this integral is numerically delicate due to singularities at $\omega=0$ and $\omega^{\prime}=0$. To remove these singularities, we introduce the logarithmic change of variables $\omega=e^{s}$, which yields
$\tilde{\psi}(s)=-\frac{1}{2}\int_{\ln(\pi T)}^{\ln(\omega_{c})}\big[U(e^{s}-e^{s^{\prime}})+U(e^{s}+e^{s^{\prime}})\big]\tilde{\psi}(s^{\prime})\,ds^{\prime},$ (45)
where $\tilde{\psi}(s)=e^{s/2}\psi(e^{s})$. This transformation renders the kernel smooth and non-singular, allowing for stable iterative numerical solution.
The resulting $T_{c}$ for the three dimensional electron gas at $r_{s}=1$ is shown in Fig. 10 (a). The $x$ and $y$ axes of the figure represent the two cut-offs $\log_{10}(k_{c}/k_{F})$ and $\omega_{c}/E_{F}$. As discussed at the beginning of this section, the momentum cut-off $k_{c}$ is a limit on the size of the superconducting gap. Fig. 10 (a) suggests that it is possible to increase $T_{c}$ at lower $\omega_{c}$ simply by going to smaller $k_{c}$, which is not a physical cut-off in the problem. However, this is actually a limitation of using the $T=0$ screened Coulomb interaction for our problem, which does not smear the momentum singularities. Therefore, the interaction enhancement below $k_{c}\gtrsim T_{c}/v_{F}$ are artifacts of the zero temperature approximation for the RPA screened Coulomb interaction. The other cut-off $\omega_{c}$, as also discussed in the beginning of a section is, unfortunately, a phenomenological parameter in the theory of SC and represents the unknown cut-off beyond which higher loop corrections must be included into Eliashberg theory. Since this is an unknown we plot the results in Fig. 10 (a) as a function of $\omega_{c}$. Consistent with the previous sub-sections, there is no significant superconductivity below $\omega_{c}<0.5E_{F}$. $T_{c}$ for the three-dimensional electron gas at $r_{s}=5$, shown in Fig.10 (b) is qualitatively similar to the $r_{s}=1$ case. The main difference between the two resuls are at large momentum cut-off $k_{c}$, where $T_{c}$ is significantly enhanced compared to the $r_{s}=1$ case. Thus, the enhancement of $T_{c}$ with increasing $r_{s}$ in 3D is mostly an artifact of an unphysical choice of $k_{c}$ related to the zero-temperature approximation RPA screened Coulomb interaction.
The corresponding results for the 2D electron at $r_{s}=1$ as shown in Fig. 10 (c), are notably different. Here, variations in $T_{c}$ have a linear dependence on $k_{c}$ as opposed to the logarithmic dependence in 3D. This motivates the use of a linear $k_{c}$ axis in panels (c,d) of Fig. 10. When $\omega_{c}\sim E_{F}$ and $k_{c}$ is large, $T_{c}$ is significantly higher than in 3D. Similar to the 3D case (using the constraint $k_{c}\gtrsim T_{c}/v_{F}$ ), there is no significant superconductivity for $\omega_{c}\lesssim 0.1E_{F}$ . For $r_{s}=5$ in 2D, shown in Fig.10 (d), The variations relative to $r_{s}=1$ are small: $T_{c}$ is slightly enhanced at large $k_{c}$ but decreases at small $k_{c}$.
@@ V Summary and Conclusion
where $\omega_{p}$ is the plasma frequency. We note that Eqs. (46) and (47) here, while being the leading order results in $q$ and $\omega$ respectively, refer to the limits $qv_{F}\gg\omega$ and $qv_{F}\ll\omega$, respectively. We note that Eqs. (46) and (47) respectively indicate the low frequency screening and the high frequency anti-screening behaviors, with the screened interaction $u(q,\omega)=v_{c}(q)/\epsilon(q,\omega)$ behaving as $u<v_{c}$ and $u>v$ for $\omega\ll qv_{F}$ and $\omega\gg qv_{F}$. Note also that for $\omega\gg\omega_{p}$, the Coulomb interaction is unscreened. All of these results are exact, and do not lead to any change in the sign of the effective or screened interaction, and this cannot lead to an effective attraction arising from electron-electron interaction. We note the crucial feature that, although both the high- (Eq. (47)) and low- (Eq. (46)) effective interaction remains repulsive, there is an intermediate frequency regime defined by $qv_{F}<\omega<\omega_{p}$, where the effective dynamically screened interaction becomes attractive since in this regime of $\omega<\omega_{p}$, the dielectric screening in Eq. (47) is negative. This emergent attraction between electrons is entirely a many-body retardation effect, which could, in principle, lead to pairing. The actual attractive region, as described in Sec. II, is larger than the asymptotic considerations of Eqs. (46) and (47) indicate. In 2D, the $q^{2}/q_{TF}^{2}$ factor in Eq. (46) is replaced by $q/q_{TF}$, but this does not change any of our considerations discussed above.
@@
In Secs. III and IV, we provide the results for $T_{c}$ based on the minimal BCS theory and the more complete Migdal-Eliashberg theory, respectively. The naive BCS theory predicts absurdly high $T_{c}$ ($\sim 100$$1000$ K) for normal metals arising from plasmon-induced pairing, with the fundamental reason behind this absurdity being the very high electronic energy scales of metals ($\sim 10^{4}$ K) and the basic electron-plasmon coupling strength, $r_{s}$, also being large, $r_{s}\sim 5$ for metals. There are two related reasons for the manifest failure of the BCS theory here: the presence of the underlying omnipresent Coulomb repulsion (i.e. the so-called $\mu^{*}$ effect) suppressing $T_{c}$ and the vertex corrections. Neither of these effects is included in the BCS theory, making its predictions unreliable. An ad hoc inclusion of a $\mu^{*}$ could certainly push $T_{c}$ down, and in fact it is easy to suppress $T_{c}$ to zero simply by positing that $\mu^{*}$ is very large in this problem—e.g. a $\mu^{*}>1$ will endure that normal metals do not undergo plasmon induced superconductivity, but a hypothetical system with very low density (and very large $r_{s}$) still might since such a large $r_{s}$ enables overcoming the repulsive $\mu^{*}$ effect. This is, however, not doing theoretical physics, it is just data fitting, since there is no reliable method for calculating $\mu^{*}$ quantitatively. The problem of vertex corrections is conceptually even greater. The quantitative success of the theory of superconductivity for electron-phonon interaction is based entirely on the existence of a small parameter (even when the basic coupling is strong, as, e.g., for Pb where the dimensionless electron-phonon coupling $\sim 1.7$) which is essentially the ratio of the bosonic energy scale to the fermionic energy scale (or their group velocity ratio). For phonons, this ratio is very small since the phonon Debye energy $\sim 10^{2}$ K and the electron Fermi energy $\sim 10^{4}$ K. This is the celebrated Migdal theorem which ensures that vertex corrections are negligible, leading to the Migdal-Eliashberg theory. There is no Migdal theorem whatsoever for the electron-electron interaction induced superconductivity by virtue of the fact that both the bosonic glue (e.g. plasmons) and the electrons, by definition, have the same energy scales, and hence the energy (or velocity) scale difference, which is the key to the Migdal theorem, does not exist for this problem by definition. The nonexistence of the Migdal theorem makes any calculation of $T_{c}$ inherently unreliable for the current problem. In Sec. IV, we elaborate on this matter by providing a detailed analysis of the Migdal-Eliashberg theory as applied to 3D normal metals in the context of the electron interaction induced superconductivity. The application of Eliashberg theory, because of Migdal’s theorem, requires us to identify a frequency cut-off $\omega_{c}$, where the Eliashberg gap equation is solved. Additionally, the $T=0$ approximation used for screening required us to consider a finite momentum space cutoff as well. For this analysis we effectively include both real and imaginary parts of the RPA dielectric function by using imaginary frequencies, so that pairing from both plasmons and electron-hole pairs are included. Because the solution of the Eliashberg equation at very small $T_{c}$ can be subtle, we derive a condition for existence of superconductivity together with an estimate for $T_{c}$. Choosing reasonable values of the cut-off, $\omega_{c}$ we find that the electron-gas to be non-superconducting in much of the range of $r_{s}$. However, by calculating $T_{c}$ from numerically solving the Eliashberg equation in Sec. IV.5 we find that a large cutoff $\omega_{c}\sim E_{F}$ can lead to an unrealistically large value of $T_{c}$. Corrections to RPA, which is the basis of our interaction model, are likely to become more important at higher $r_{s}$. The range of $r_{s}$ over which the gas is superconducting increases with lowering $k_{c}$. However, a $k_{c}$ much smaller than $10^{-3}k_{F}$ likely corresponds to a sub-kelvin gap superconductor. Furthermore, the superconducting gap or even the presence of superconductivity is found to be drastically reduced by lowering $\omega_{c}$, which is likely required by Migdall’s theorem to avoid significant vertex correction effects. Ultimately, $\omega_{c}$ needs to be chosen to be small to avoid vertex corrections but too low a value of $\omega_{c}$ would under-estimate $T_{c}$. Our conclusion is that in all likelihood any plasmon induced $s$-wave metallic superconductivity has vanishing $T_{c}$. In fact, we find that even the $T=0$ system is not ordered for small eneough $\omega_{c}$, i.e., electron-electron interaction simply does not lead to $s$-wave superconductivity in 3D normal metals. It is still however possible that the approximations in our analysis of the Eliashberg equation miss a SC state with a very low $T_{c}$. We mention that our work does not rule out an exponentially low $T_{c}$ for normal metals in higher angular momentum channels arising from the Kohn-Luttinger mechanism, which we do not consider in the current work.
@@
One could legitimately ask whether our conclusion of there being no plasmon-induced superconductivity applies to systems other than 3D jellium electron liquids, which is the explicit case we consider. We believe that our work should apply to all situations where a jellium electron gas model applies since there is nothing specific to a 3D model of metals utilized in our work. The fact that energy scales cannot be separated, and hence vertex corrections and Coulomb repulsion effects are important, but cannot be included in the BCS-Migdal-Eliashberg theories, applies to all situations, including much-studied 2D electron liquids, where the effective $r_{s}$ is often larger (and the effective $\omega_{p}$ smaller) because of 2D electron liquids, mostly being doped semiconductors, are inherently low-density metals. The fact that 2D plasmon dispersion goes as $q^{1/2}$ and thus vanishes at long wavelength is a matter of profound inconsequence in our theoretical considerations since all the results involve some type of frequency-momentum integrations, and what matters is the plasmon energy at $k_{F}$ (or the plasmon energy at the momentum where it enters the electron-hole continuum). In fact, the 2D theory for plasmon-induced metallic superconductivity is identical to our 3D theory, with exactly the same conclusions. There have been claims [Takada1978] that having a bilayer 2D systems with acoustic plasmons [DasSarma1982] may facilitate plasmon-induced superconductivity, perhaps motivated by the fact that acoustic plasmons have the same linear-in-momentum energy dispersion as acoustic phonons. But this claim is misguided and misleading because the Migdal theorem still does not apply, and the standard SC theories (i.e. BCS-Migdal-Eliashberg) would produce results very similar to what we get in our work, again with unreliable and unreasonably high $T_{c}$. The problem of the inapplicability of the Migdal theorem does not become any less stringent just by virtue of having acoustic plasmons in the system since the bosons and the fermions are all arising from exactly the same Hamiltonian. As described in Sec. IV.4, the Eliashberg-Migdall analysis described in the previous paragraph can be extended to 2D electron gases as well. The results are consistent with the argument in this paragraph, despite the lower plasma frequency in 2D and the superconducting $T_{c}$ for the 2D electron gas is vanishingly small at $\omega_{c}\lesssim 0.1E_{F}$. However, the values of cutoff where $T_{c}$ is suppressed needs to be chosen lower in the 2D electron gas compared to the 3D case.
@@
We mention that the manifest failure of the Migdal theorem for a pairing mechanism does not immediately lead to the conclusion that superconductivity may not occur due to that mechanism. All it implies is the failure of our existing theories, the BCS-Migdal-Eliashberg theory, to describe any possible superconductivity arising from that mechanism. Thus, the absurdity of blindly applying the BCS theory for plasmon induced pairing in 3D metals leading to a predicted $T_{c}\sim 10^{2}$$10^{3}$ K only implies, by itself, that such a theory is not useful in discussing superconductivity in this context. For example, if the problem could be solved exactly in some magical manner, where the issue of the inapplicability of the Migdal theorem is no longer relevant, we cannot rule out the existence of a plasmon-induced $T_{c}$. The work in our Sec. IV indicates such a scenario to be extremely unlikely, but we have not solved the problem exactly. We do believe that if the plasmon mechanism, in an unlikely (but not impossible) scenario, leads to superconductivity in normal metals, the associated $T_{c}$ is likely to be impractically small by virtue of the fact that the direct effect of Coulomb repulsion will always be a huge challenge, i.e., the effective mu* is likely to be large negating any attractive pairing trend induced by plasmons. The fact that the bare Coulomb interaction between electrons is repulsive added to the fact that plasmons are intrinsic electronic phenomena make it essentially impossible for the plasmon mechanism to lead to any ‘practical’ conventional superconductivity in our opinion.
@@
We now comment on the possibility of electron interaction induced superconductivity in strictly lattice strongly correlated Hamiltonians described by a tight binding Hamiltonian with mostly local interactions. The paradigmatic model is the Hubbard model with onsite interaction and nearest neighbor hopping. This model, which is complementary to our jellium based free electron Hamiltonian with long range Coulomb interaction, has been much studied in the literature motivated by high-$T_{c}$ cuprate superconductors. Our work does not directly apply to this model although the inapplicability of the Migdal theorem as well as the $\mu^{*}$ problem exist here too. The question of superconductivity in the repulsive Hubbard model is an extensively studied problem, which is well beyond the scope of the current work. We mention two relevant works. In Ref. [Raghu2010], the authors conclude that for asymptotically weak interaction, where the weak coupling perturbative renormalization group should apply, there is a $T_{c}$ which is exponentially small. This is akin to the celebrated Kohn-Luttinger superconductivity in electron liquids. The authors speculate that if their perturbative results can be uncritically extended to larger interaction strength, then a $T_{c}$ describing ‘regular’ superconductivity (i.e. not exponentially small) may emerge. But, this conclusion has been directly refuted in a later quantum Monte Carlo and DMRG calculations which assertively claims the absence of superconductivity in the repulsive Hubbard model for any interaction strength [Qin2020]. But a later numerical work found that a slight modification of the tight binding kinetic energy term to include the next-nearest-neighbor hopping in a generalized Hubbard model does manifest superconductivity, but most likely with rather low $T_{c}$, and in higher angular momentum channel (e.g. $d$-wave) [Xu2024]. We believe that the superconductivity discussed in the context of the Hubbard model in these (and possibly other) publications are essentially low-temperature superconductivity in higher orbital angular momentum, whose physics is connected to Kohn-Luttinger superconductivity in the electron liquid context, and not the ‘conventional’ superconductivity with ’practical’ transition temperature of interest in the current work. Certainly, the superconductivity discussed in the strongly correlated lattice models of electron interactions is not plasmon-induced in the sense we describe in the current work, and is therefore outside the scope of the current work. In this context, there are recent theories of superconductivity in various graphene and TMD based 2D layers with rather flatbands (which are very far from our electron liquid metals in the current work), where the superconductivity (with low $T_{c}$) may very well be arising from electron-electron interactions, but again such superconductivity is most likely a variation on the Kohn-Luttinger mechanism and is not generated by the attractive pairing glue of the virtual plasmon exchange between the electrons [Chou2021, Chou2025, Zhu2025].